a solution has a hydroxide-ion concentration of 1.0 x 10^-3 mol per liter. what is the pH of this solution?
Answer #1

hydroxide ion concentration, [OH-] = 1.0 x 10-3 mol/L

pOH = -log[OH-]

pOH = -log(1.0 x 10-3 mol/L)

pOH = 3.0

pH = 14 - pOH

pH = 14 - 3.0

pH = 11.0